OpenStudy (anonymous):

(ALGEBRA) If px2 - mx - c = 0, what is x equal to?

OpenStudy (anonymous):

if \[ax^2+bx+c=0\] then \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

put \(a=p,b=-m,c=-c\)

OpenStudy (anonymous):

if it is not clear what the answer is, let me know and i will write it

OpenStudy (anonymous):

I am confused by all of this subject matter

OpenStudy (anonymous):

it is a matter of substituting one set of variables for another there is no content here at all

OpenStudy (anonymous):

\[\huge x=\frac{-\color{blue}b\pm\sqrt{\color{blue}b^2-4\color{red}a\color{green}c}}{2\color{red}a}\]

OpenStudy (anonymous):

in this case \[\huge \color{red}a=\color{red}p, \color{blue}b=\color{blue}{-m}, \color{green}c=\color{green}{-c}\]

OpenStudy (anonymous):

when you substitute you get \[\large x=\frac{m\pm\sqrt{m^2+4pc}}{2p}\]